OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)? A-(x − 2)2 + (y + 3)2 = 5 B-(x + 2)2 + (y − 3)2 = 25 C-(x − 2)2 + (y + 3)2 = 25 D-(x − 2)2 − (y + 3)2 = 5

Nnesha (nnesha):

equation of circle is \[\huge\rm (x-h)^2 + (y -k)^2 = r^2\]where (h,k) is the center r = radius and (-2,0) = (x,y)

Nnesha (nnesha):

replace variables by their values solve for r

OpenStudy (anonymous):

Thank you!

Nnesha (nnesha):

my pleasure :=)